OpenStudy (anonymous):

how do you factor 9a^2+4bc-4c^2-b^2? PLEASE?!?!?!?!??!

OpenStudy (anonymous):

Not much you can do with that..

OpenStudy (anonymous):

um. yea. thanks

OpenStudy (anonymous):

There is some partial factoring you can do if you separate out some differences of squares, e.g.

OpenStudy (anonymous):

Try this and see where it goes: \[9a^2-(b^2-4bc+4c^2)\]

OpenStudy (anonymous):

i did that, and i got 3a-(b^2-4bc+4c^2) and now i have no idea what to do

OpenStudy (anonymous):

\[\rightarrow 9a^2-(b-2c)^2\]

OpenStudy (anonymous):

What happened to the other 3a in 9a^2?

OpenStudy (anonymous):

OH!!!!!! so you turn it into a correct trinomial by changing the addition sign to a subtraction sign in front of the parenthesis, changing the signs of all of the terms! then you factor that correct trinomial to get (3a-b-2c)(3a-b-2c)

OpenStudy (anonymous):

+2c at the end, otherwise, bingo!

OpenStudy (anonymous):

oh. thanksss!!!!!!!!!! :D

OpenStudy (anonymous):

I retract my earlier statement about there not being much to do; took me a minute to see all those squares.

OpenStudy (anonymous):

haha ok(:

OpenStudy (anonymous):

(3a+b-2c)(3a-b-2c)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Oops, (3a+b-2c)(3a-b+2c)

OpenStudy (anonymous):

? Did I not distribute the negative sign correctly?

OpenStudy (anonymous):

Yeah, going from here: (3a+(b-2c))(3a-(b-2c)) you should get (3a+b-2c)(3a-b+2c)

OpenStudy (anonymous):

oh, yes, and +b at the beginning. That's what I got too. I didn't catch that error in Hannah's

OpenStudy (anonymous):

oh yes @jabberwock hes right, because a -b times a -b equals positive b^2

OpenStudy (anonymous):

But you're using x^2-y^2=(x+y)(x-y)

OpenStudy (anonymous):

wait, now im confusedddd..........so the original problem is 9a+(4bc-4c^2-b^2) right? then i changed the positive in front of the parenthesis to a negative. i get that. but i dont understand why you would change that. but we did anyways. so then i got 9a-(b^2-4bc+4c^2) because all the signs would be opposite. then we factored that and got (3a+b-2c)(3a-b+2c). I have no idea why. care to explain??

OpenStudy (anonymous):

(3a+b-2c)(3a-b+2c) is correct. \[\rightarrow 9a^2-3ab+6ac+3ab-b^2+2bc-6ac+2bc-4c^2\]

OpenStudy (anonymous):

\[9a^2-(b^2-4bc+4c^2)\]\[9a^2-(b-2c)^2\] Good to there?

OpenStudy (anonymous):

@jabberwock yes, but i dont understand why you would change that positive in front of the parenthesis to a negative

OpenStudy (anonymous):

Isn't that how it appeared in the original problem?

OpenStudy (anonymous):

Oh...it's not written that way in what you wrote...

OpenStudy (anonymous):

nope! in the original it was 9aPLUS(4bc-4c^2-b^2)

OpenStudy (anonymous):

so why would you change that plus to a minus?

OpenStudy (anonymous):

Ah, you're right. CliffSedge and I were both thinking it was (-)... \[9a^2+(b-2c)^2\]is as far as you can go.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

OH, nevermind again. Heh, sorry. Original problem: \[9a^2+4bc-4c^2-b^2\] Rearrange the terms: \[9a^2-b^2+4bc-4c^2\]

OpenStudy (anonymous):

ok,

OpenStudy (anonymous):

Notice: \[-b^2+4bc-4c^2\]is the same as\[-(b^2-4bc+4c^2)\]by factoring out -1

OpenStudy (anonymous):

ok,

OpenStudy (anonymous):

oh, so you just put a parenthesis around the trinomial, and distributed the -1 and changed all of the signs.!

OpenStudy (anonymous):

Exactly :)

OpenStudy (anonymous):

ok then you factor to (3a-b+2c)(3a-b+2c)!!!~

OpenStudy (anonymous):

Right :) :)

OpenStudy (anonymous):

No, wait

OpenStudy (anonymous):

Not right

OpenStudy (anonymous):

thanks so much!!!!!!!---what?? lol

OpenStudy (anonymous):

\[9a^2-(b-2c)^2\] Notice that you have two squared terms. You have \[(3a)^2-(b-2c)^2\] Yes?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

So this is called the "difference of two squares," which happens whenever you have a squared term minus another squared term. It ALWAYS factors like this: \[m^2-n^2=(m+n)(m-n)\]

OpenStudy (anonymous):

yes i learned that,

OpenStudy (anonymous):

K, so in this case m = 3a n = 2b-c What does it look like when you substitute?

OpenStudy (anonymous):

(3a+2b-c)(3a-2b-c)

OpenStudy (anonymous):

don't you mean n=b-2c?

OpenStudy (anonymous):

Oops for like the 4th time...lol, sorry Almost. \[(3a+(b-2c))(3a-(b-2c))\] Look at the -(b - 2c) and distribute the negative to get -b + 2c

OpenStudy (anonymous):

OK!!!!!! haha lol(: thankkkssss:D

OpenStudy (anonymous):

Sure, sorry about all the mistakes.

OpenStudy (anonymous):

no prob(: